OpenStudy (anonymous):

I have tried this like 6 times: please show how problem is solved.Jill Corporation accepted a $16,000 note on Aug. 12. Terms of the note were 13% for 100 days. Jill discounted the note on September 28 at the Reno Bank at 14%. The proceeds to Jill would be:

OpenStudy (nottim):

Man, you are going to have to define some things.

OpenStudy (anonymous):

is a note money? 13% for 100 days- an increase in what she got or...? discounted? yeah I am really sorry but I'm not sure what you are asking

OpenStudy (anonymous):

ok watched a video and here is what I think this answer is: Date of Note Face Value Length Interest Bank Discount Date of of note of note rate rate discount Aug 12. $16,000 100 days .13 .14 Sept 28 Interest = $16,000 * .13 * 100/360 = $577.78 16000-577.78 = 15422.22 Aug 12 + 100 days = Nov 20 Sept 28 to Nov 20 is 53 days 16000 * .14 * 53/360 = 2239.85 16000 – 2239.85 = 13760.15 (proceeds)